OpenStudy (anonymous):

What is the 24th term of the arithmetic sequence where a1 = 8 and a9 = 56?

OpenStudy (campbell_st):

all you need to common difference. the easy way is \[d =\frac{ (56 - 8)}{(9 - 1) }\] when you get that you'll need to formula for a term in an arithmetic sequence \[a_{N} = a_{1} + ( n - 1) \times d\] you know n = 24 and you know a1 = 8 hope it helps

OpenStudy (anonymous):

d = 6 an= 8 + (24-1) x 6

OpenStudy (campbell_st):

thst looks good to mr

OpenStudy (anonymous):

an = 8 + 23 x 6 = 8 + 138 = 146?

OpenStudy (anonymous):

Thank you very much